if A=30 degree , a=7,b=8,in triangle ABC, then prove that b has two solutions

Grade:11

1 Answers

sahil
6 years ago
se this question is on ambiguity solution:by sine rule Sin(a)/a=Sin(b)/bSin(b)=b×sin(a)/a=8×sin(30)/7=4/7 since sin(b) is less than unity it has two solution one between 0 to 90 and other between 90 to 180. I tried to give ans in simplest way if you are satisfied please click on approve.